OpenStudy (anonymous):

Evaluate P(3, 0). 0 1 6

OpenStudy (blockcolder):

That's like saying, "In how many ways can 3 people sit in 0 chairs?" And of course, there's only one way: where they don't sit at all.

OpenStudy (anonymous):

0

OpenStudy (blockcolder):

Using the formula gives the same result: \[\frac{3!}{(3-0)!}=\frac{3!}{3!}=1\]